OpenStudy (anonymous):

Can someone please help me solve for "x". 2cos^2 x = sec^2 x / 2

OpenStudy (mathmate):

hint: substitute \(\sec(x) = \dfrac{1}{\cos(x)}\).

OpenStudy (mathmate):

Is question: A. \(2\cos^2(x)=\dfrac{\sec^2(x)}{2}\) or B. \(2\cos^2(x)=\sec^2(\dfrac{x}{2})\)